METAL ION CHEMISTRY IN COLD INTERSTELLAR AND CIRCUMSTELLAR ENVIRONMENTS

\underline{ROBERT C. DUNBAR}, Chemistry Department, Case Western Reserve University, Cleveland, OH 44106; and SIMON PETRIE, Chemistry Department, the Faculties, Australian National University, Canberra, ACT, Australia.

Ions of abundant metals, including Na^{+}, K^{+}, Mg^{+}, Ca^{+} and Fe^{+}, are likely to play an important role in the chemical evolution of metal-containing interstellar and circumstellar environments. The ionic reaction networks include a variety of efficient radiative association, ligand-switching, clustering and recombination processes. High-level ab initio calculations, using the CP-dG2thaw method and related approaches, are combined here with VTST computations of reaction kinetics to explore the reactivity of the atomic metal ions and metal-containing molecular ions with abundant neutrals. The likely composition of the metal-containing populations and the accompanying formation of neutral species like MNC, MCN and MC_{3}N are examined for three different C-rich environments, typified by the cold dark interstellar cloud TMC-1, the outer circumstellar envelope of the AGB star IRC+10216 and the post-AGB object CRL 618, and are discussed with reference to molecular detections within these three objects.